OpenStudy (anonymous):

simplify 5 ^3sqrt 163 135 ^3sqrt 6 15 ^3sqrt 6 8 ^3sqrt 6 3 ^3sqrt 6

OpenStudy (anonymous):

By ^3sqrt 163 do you mean\[\large \sqrt[3]{163}\]

OpenStudy (anonymous):

my bad it is \[5\sqrt[3]{162}\]

OpenStudy (anonymous):

162 or 163?

OpenStudy (anonymous):

162, made a typo

OpenStudy (anonymous):

3 * 3 * 3 * 3 * 2 Answer is 15 root 6

OpenStudy (anonymous):

cube root, sorry

OpenStudy (anonymous):

\[162 = 81\cdot 2 = 3\cdot 3\cdot 3\cdot 3\cdot 2 =3^3 \cdot 3 \cdot 2 \]\[\implies 5\sqrt[3]{162} = 5\sqrt[3]{3^3\cdot 3\cdot 2}=5(3)\sqrt[3]{3\cdot 2}\]